OpenStudy (anonymous):

Write an appropriate direct variation equation if y = 8 when x = 2. *metals*

OpenStudy (campbell_st):

well the basic direct variation equation is y = kx k is the constant of variation so substituting \[8 = k \times 2\] what do you think the value of k is?

OpenStudy (anonymous):

y =16/x y = 4x y = 8x y =4/x (options)

OpenStudy (anonymous):

I think it's y=4x

OpenStudy (campbell_st):

well done

OpenStudy (campbell_st):

are you using the same direct variation equation...?

OpenStudy (anonymous):

I think so but it's graphing

OpenStudy (campbell_st):

well find the slope of the line.... and the slope will be the k value

OpenStudy (anonymous):

I think it's direct

OpenStudy (campbell_st):

opps... misread the equation.... because its a straight line its a direct variation the value of k... is the slope of the line

OpenStudy (anonymous):

so it's inverse

OpenStudy (campbell_st):

no, its direct..... y = 1/3 x

OpenStudy (anonymous):

If y varies inversely as x, and y = 4 as x = 8, find y for the x-value of 2. (next work)